\begin{thebibliography}{}
\expandafter\ifx\csname natexlab\endcsname\relax\def\natexlab#1{#1}\fi

\bibitem[{Black et~al.(2011)Black, Burger, Conkie, Hastie, Keizer, Lemon,
  Merigaud, Parent, Schubiner, Thomson, Williams, Yu, Young, and
  Eskenazi}]{black-EtAl:2011:SIGDIAL2011}
Alan~W Black, Susanne Burger, Alistair Conkie, Helen Hastie, Simon Keizer,
  Oliver Lemon, Nicolas Merigaud, Gabriel Parent, Gabriel Schubiner, Blaise
  Thomson, Jason~D. Williams, Kai Yu, Steve Young, and Maxine Eskenazi. 2011.
\newblock Spoken dialog
  challenge 2010: Comparison of live and control test results.
\newblock In {\em Proceedings of the SIGDIAL 2011 Conference\/}. Association
  for Computational Linguistics, Portland, Oregon, pages 2--7.
\newblock
  \href{http://www.aclweb.org/anthology/W11-2002}{http://www.aclweb.org/anthology/W11-2002}.

\bibitem[{Du\v{s}ek and Jur\v{c}{\'{\i}}\v{c}ek(2015)}]{jurcicek:2015:ACL}
Ond\v{r}ej Du\v{s}ek and Filip Jur\v{c}{\'{\i}}\v{c}ek. 2015.
\newblock Training a natural
  language generator from unaligned data.
\newblock In {\em Proc.\ of ACL-IJCNLP (Volume~1: Long Papers)\/}. Beijing,
  China, pages 451--461.
\newblock
  \href{http://aclweb.org/anthology/P15-1044}{http://aclweb.org/anthology/P15-1044}.

\bibitem[{Du\v{s}ek and Jur\v{c}{\'{\i}}\v{c}ek(2016)}]{Dusek:ACL16}
Ond\v{r}ej Du\v{s}ek and Filip Jur\v{c}{\'{\i}}\v{c}ek. 2016.
\newblock Sequence-to-sequence
  generation for spoken dialogue via deep syntax trees and strings.
\newblock In {\em Proc.\ of ACL\/}. Berlin, Germany, pages 45--51.
\newblock {arXiv:1606.05491}.
\newblock
  \href{http://aclweb.org/anthology/P16-2008}{http://aclweb.org/anthology/P16-2008}.

\bibitem[{Gkatzia and Mahamood(2015)}]{gkatzia:enlg2015}
Dimitra Gkatzia and Saad Mahamood. 2015.
\newblock A snapshot of {NLG}
  evaluation practices 2005--2014.
\newblock In {\em Proc.\ of ENLG\/}. Brighton, UK, pages 57--60.
\newblock
  \href{https://doi.org/10.18653/v1/W15-4708}{https://doi.org/10.18653/v1/W15-4708}.

\bibitem[{Han et~al.(2013)Han, Kashyap, Finin, Mayfield, and
  Weese}]{han2013umbc}
Lushan Han, Abhay Kashyap, Tim Finin, James Mayfield, and Jonathan Weese. 2013.
\newblock {UMBC\_EBIQUITY-CORE}:
  Semantic textual similarity systems.
\newblock In {\em Proc.\ of *SEM\/}. Atlanta, Georgia, volume~1, pages 44--52.
\newblock
  \href{http://aclweb.org/anthology/S13-1005}{http://aclweb.org/anthology/S13-1005}.

\bibitem[{Lampouras and Vlachos(2016)}]{vlachos:coling2016}
Gerasimos Lampouras and Andreas Vlachos. 2016.
\newblock Imitation learning for
  language generation from unaligned data.
\newblock In {\em Proc.\ of COLING: Technical Papers\/}. Osaka, Japan, pages
  1101--1112.
\newblock
  \href{http://aclweb.org/anthology/C16-1105}{http://aclweb.org/anthology/C16-1105}.

\bibitem[{Lin(2004)}]{lin2004rouge}
Chin-Yew Lin. 2004.
\newblock {ROUGE}: A package for
  automatic evaluation of summaries.
\newblock In {\em Text summarization branches out: Proc.\ of the ACL-04
  workshop\/}. Barcelona, Spain, pages 74--81.
\newblock
  \href{http://aclweb.org/anthology/W04-1013}{http://aclweb.org/anthology/W04-1013}.

\bibitem[{Lu(2009)}]{lu2009automatic}
Xiaofei Lu. 2009.
\newblock Automatic measurement of
  syntactic complexity in child language acquisition.
\newblock {\em International Journal of Corpus Linguistics\/} 14(1):3--28.
\newblock
  \href{http://doi.org/10.1075/ijcl.14.1.02lu}{http://doi.org/10.1075/ijcl.14.1.02lu}.

\bibitem[{Lu(2012)}]{lu2012relationship}
Xiaofei Lu. 2012.
\newblock The
  relationship of lexical richness to the quality of esl learners’ oral
  narratives.
\newblock {\em The Modern Language Journal\/} 96(2):190--208.
\newblock
  \href{http://doi.org/10.1111/j.1540-4781.2011.01232\_1.x}{http://doi.org/10.1111/j.1540-4781.2011.01232\_1.x}.

\bibitem[{Mairesse et~al.(2010)Mairesse, Ga{\v{s}}i{\'c},
  Jur{\v{c}}{\'\i}{\v{c}}ek, Keizer, Thomson, Yu, and Young}]{mairesse:acl2010}
Fran{\c{c}}ois Mairesse, Milica Ga{\v{s}}i{\'c}, Filip
  Jur{\v{c}}{\'\i}{\v{c}}ek, Simon Keizer, Blaise Thomson, Kai Yu, and Steve
  Young. 2010.
\newblock Phrase-based statistical
  language generation using graphical models and active learning.
\newblock In {\em Proc.\ of ACL\/}. Uppsala, Sweden, pages 1552--1561.
\newblock
  \href{http://aclweb.org/anthology/P10-1157}{http://aclweb.org/anthology/P10-1157}.

\bibitem[{Mei et~al.(2016)Mei, Bansal, and Walter}]{Mei:NAACL2015}
Hongyuan Mei, Mohit Bansal, and Matthew~R. Walter. 2016.
\newblock What to talk about and
  how? {Selective} generation using {LSTMs} with coarse-to-fine alignment.
\newblock In {\em Proc.\ of NAACL-HLT 2016\/}. San Diego, CA, USA.
\newblock {arXiv:1509.00838}.
\newblock
  \href{http://aclweb.org/anthology/N16-1086}{http://aclweb.org/anthology/N16-1086}.

\bibitem[{Novikova and Rieser(2016)}]{novikova2016crowd}
Jekaterina Novikova and Verena Rieser. 2016.
\newblock
  Crowd-sourcing
  {NLG} data: Pictures elicit better data.
\newblock In {\em Proc. of the 9th International Natural Language Generation
  conference (INLG)\/}.
\newblock
  http://www.aclweb.org/anthology/W/W16/W16-6644.pdf.

\bibitem[{Papineni et~al.(2002)Papineni, Roukos, Ward, and
  Zhu}]{papineni2002bleu}
Kishore Papineni, Salim Roukos, Todd Ward, and Wei-Jing Zhu. 2002.
\newblock {BLEU}: a method for
  automatic evaluation of machine translation.
\newblock In {\em Proc.\ of ACL\/}. Philadelphia, PA, USA, pages 311--318.
\newblock
  \href{http://aclweb.org/anthology/P02-1040}{http://aclweb.org/anthology/P02-1040}.

\bibitem[{Snover et~al.(2006)Snover, Dorr, Schwartz, Micciulla, and
  Makhoul}]{ter}
Matthew Snover, Bonnie Dorr, Richard Schwartz, Linnea Micciulla, and John
  Makhoul. 2006.
\newblock A study of
  translation edit rate with targeted human annotation.
\newblock In {\em Proc.\ of AMTA\/}. Cambridge, MA, USA, pages 223--231.
\newblock
  \href{http://mt-archive.info/AMTA-2006-Snover.pdf}{http://mt-archive.info/AMTA-2006-Snover.pdf}.

\bibitem[{Specia et~al.(2010)Specia, Raj, and Turchi}]{specia:MT2010}
Lucia Specia, Dhwaj Raj, and Marco Turchi. 2010.
\newblock Machine translation
  evaluation versus quality estimation.
\newblock {\em Machine translation\/} 24(1):39--50.
\newblock
  \href{https://doi.org/10.1007/s10590-010-9077-2}{https://doi.org/10.1007/s10590-010-9077-2}.

\bibitem[{Wen et~al.(2016)Wen, Ga\v{s}i\'c, Mrk\v{s}i\'c, Rojas{-}Barahona, Su,
  Vandyke, and Young}]{Wen:NAACL16}
Tsung{-}Hsien Wen, Milica Ga\v{s}i\'c, Nikola Mrk\v{s}i\'c, Lina~Maria
  Rojas{-}Barahona, Pei{-}Hao Su, David Vandyke, and Steve~J. Young. 2016.
\newblock Multi-domain neural
  network language generation for spoken dialogue systems.
\newblock In {\em Proc.\ of EMNLP-HLT\/}. San Diego, CA, USA, pages 120--129.
\newblock {arXiv:1603.01232}.
\newblock
  \href{http://aclweb.org/anthology/N16-1015}{http://aclweb.org/anthology/N16-1015}.

\bibitem[{Wen et~al.(2015)Wen, Ga\v{s}i\'c, Mrk\v{s}i\'c, Su, Vandyke, and
  Young}]{wen:emnlp2015}
Tsung-Hsien Wen, Milica Ga\v{s}i\'c, Nikola Mrk\v{s}i\'c, Pei-Hao Su, David
  Vandyke, and Steve Young. 2015.
\newblock Semantically conditioned
  {LSTM}-based natural language generation for spoken dialogue systems.
\newblock In {\em Proc.\ of EMNLP\/}. Lisbon, Portugal, pages 1711--1721.
\newblock
  \href{http://aclweb.org/anthology/D15-1199}{http://aclweb.org/anthology/D15-1199}.

\bibitem[{Williams and Young(2007)}]{Williams:2007}
Jason~D. Williams and Steve Young. 2007.
\newblock Partially
  Observable Markov Decision Processes for Spoken Dialog Systems.
\newblock {\em Comput. Speech Lang.\/} 21(2):393--422.
\newblock
  \href{https://doi.org/10.1016/j.csl.2006.06.008}{https://doi.org/10.1016/j.csl.2006.06.008}.

\end{thebibliography}

\documentclass[11pt,a4paper]{article}
\usepackage{acl2017}
\usepackage{times}
\usepackage{latexsym}
\usepackage[english]{babel}
\usepackage{blindtext}
\usepackage{adjustbox}

\usepackage{url}

\newcommand{\bleu}{\textsc{\small bleu}}
\newcommand{\rouge}{\textsc{\small rouge}}

\newcommand{\ter}{\textsc{\small ter}}

\aclfinalcopy 

\usepackage[normalem]{ulem}
\definecolor{darkgreen}{rgb}{0.0, 0.5, 0.0}

\makeatletter
\def\ODhl#1{\bgroup \markoverwith{\lower3.5\p@\hbox{\sixly \textcolor{darkgreen}{\char58}}}\ULon{#1}}
\font\sixly=lasyb10 scaled 652 
\makeatother
\def\ODdel#1{\bgroup\markoverwith{\textcolor{darkgreen}{\rule[0.5ex]{2pt}{1pt}}}\ULon{#1}}

\title{Extended Abstract for WiNLP}
\title{Addressing Bottlenecks for Data-driven Natural Language Generation}
\title{Future Directions for Data-driven Natural Language Generation}
\title{Data-driven Natural Language Generation: Paving the Road to Success}

\author{Jekaterina Novikova, Ond\v{r}ej Du\v{s}ek \and Verena Rieser\\
Interaction Lab \\
Heriot-Watt University \\
Edinburgh, EH14 4AS, UK \\
  {\tt \{j.novikova, o.dusek, v.t.rieser\}@hw.ac.uk}}

\date{}

\begin{document}
\maketitle
\begin{abstract}

We argue that there are currently two major bottlenecks 
 to the commercial use of statistical machine learning approaches for natural language generation (NLG): (a) The lack of reliable automatic evaluation metrics for NLG, and (b) The scarcity of high quality in-domain corpora. We address the first problem by thoroughly analysing current evaluation metrics and motivating the need for a new, more reliable metric. 
 The second problem is addressed  by presenting a novel framework for developing and evaluating a high quality corpus for NLG training. 
\end{abstract}

\section{Evaluation metrics for NLG}
\label{sec:metrics}

Up to 60\% of  NLG research published between 2012--2015 relies on automatic evaluation measures, such as \bleu{} \cite{gkatzia:enlg2015}. The use of such metrics is, however, only sensible if they are known to be sufficiently correlated with human preferences, which is not the case, as we show 
  in the most complete study to date, across metrics, systems, datasets and domains. 

We evaluate three end-to-end NLG systems: RNNLG \cite{wen:emnlp2015}, TGen \cite{jurcicek:2015:ACL} and LOLS \cite{vlachos:coling2016}, using a large number of 21 automated metrics. 
The metrics are divided into groups of word-based metrics (WBMs, such as {\bf \ter} \cite{ter},
{\bf \bleu} \cite{papineni2002bleu},
{\bf \rouge} \cite{lin2004rouge}, semantic similarity \cite{han2013umbc} etc.) and grammar-based metrics (GBMs, such as readability, characters per utterance and per word, syllables per sentence and per word, number of misspellings etc.). To assess the metrics' reliability, we calculate the 
Spearman correlation between the metrics and human ratings for the same natural language (NL) utterances, the accuracy of relative rankings and conduct a detailed error analysis. 

\begin{table}[tp]
\begin{center}
\begin{adjustbox}{max width=0.49\textwidth}
\begin{tabular}{l|ll|ll}
 & \multicolumn{2}{l|}{Lexical richness} & \multicolumn{2}{l}{Syntactic 
 complexity} \\
Dataset & LS & MSTTR & Level 0-1 & Level 6-7 \\
\hline \hline
our corpus & \textbf{0.57} & \textbf{0.75} & \textbf{46\%} & \textbf{16\%} \\
SFRest & 0.43 & 0.62 & 47\% & 13\% \\
SFHot & 0.43 & 0.59 & 51\% & 15\% \\
Bagel & 0.42 & 0.41 & 50\% & \textbf{16\%}
\end{tabular}
\end{adjustbox}
\end{center}
\caption{Lexical richness and syntactic variation for the collected corpus and other popular datasets. 
\textit{LS} measures the proportion of less frequent words in the text, \textit{MSTTR} measures the type-token ratio normalised by the size of the corpus. For D-level complexity, \textit{Level 0-1} include syntactically simple sentences, \textit{Level 6-7} include the most complicated sentences.}
\label{tab:res}
\end{table}

The results reveal that no metric produces an even moderate correlation with human ratings (max.\ 
$\rho=0.33$), independently of dataset, system, or aspect of human rating. WBMs make two strong assumptions: They treat human-generated NL references as a gold standard which is correct and complete. We argue that these assumptions are very often invalid for corpus-based NLG, especially when using crowdsourced datasets. Grammar-based metrics, on the other hand, do not rely on human-generated references and are not influenced by their quality. However, these metrics can be easily manipulated with grammatically correct and easily readable output that is unrelated
to the input. 
Our study clearly demonstrates the need for
more advanced metrics, as used in related fields, e.g.\ MT \cite{specia:MT2010}. 


\section{Corpus development and evaluation}
\label{sec:corpus}

Recent advances in corpus-based NLG \cite{jurcicek:2015:ACL,wen:emnlp2015,Mei:NAACL2015,Wen:NAACL16,Dusek:ACL16,vlachos:coling2016} require costly training data, consisting of meaning representations (MRs) paired with corresponding NL texts. In our work, we propose a novel framework for crowdsourcing high quality NLG training data, using automatic quality control measures and evaluating two 
types of MRs, pictorial and textual, used to elicit data \cite{novikova2016crowd}. 

When collecting corpora for training NLG systems, especially when using crowd workers, the following challenges arise:

\begin{itemize}
\item[(1)] How to ensure 
the required high quality of the collected data?
\vspace{-0.2cm}
\item[(2)] What types of meaning representations can elicit spontaneous, natural and varied data from crowd workers?
\end{itemize}

To address (1), we filter the crowdsourced data using a combination of automatic and semi-manual validation procedures, as described in \cite{novikova2016crowd}.
We validate the data by selecting native English participants, allowing only well formed English sentences to be submitted, and measuring the semantic similarity of a collected NL utterance and an associated MR. 
%
Using this framework, we collected a dataset of 50k instances in the restaurant domain, which is 10 times bigger than datasets currently used for NLG training, e.g.\ SFRest and SFHot \cite{wen:emnlp2015} 
 or Bagel \cite{mairesse:acl2010}. 

To evaluate the quality of the collected corpus, we analyse the data with regards to lexical richness and syntactic variation and compare our results to other popular datasets in similar domains, i.e.\ SFRest, SFHot and BAGEL.
We use the Lexical Complexity Analyser 
 \cite{lu2012relationship} to measure various dimensions of lexical richness and variation, such as \textit{mean segmental type-token ratio} (MSTTR) and \textit{lexical sophistication} (LS). The results of lexical analysis (see Table~\ref{tab:res}) show that our  corpus is lexically more diverse and as such, considerably more complex. 
In order to evaluate syntactic variation and complexity of NL references in our corpus, we use the D-Level Analyser \cite{lu2009automatic}. 
Table~\ref{tab:res} shows that our collected corpus has the highest proportion of  complex sentences (54\% of sentences scored above level 1). 
 At the same time, the proportion of  syntactically complex sentences (levels 6 and 7) is one of the highest. 

Furthermore, our dataset requires content selection in 40\% of the cases. In contrast to the other datasets, crowd workers were asked to verbalise all the \textit{useful} information from the MR and were allowed to skip an attribute value  considered unimportant. 
 As such, learning from this dataset promises more natural, varied and less template-like system utterances.

To address challenge (2) in corpus development, we conduct a principled study regarding the trade-off between semantic expressiveness of the MR and the quality of crowdsourced utterances elicited. 
In particular, we investigate translating textual MRs (presented in the form of logic-based dialogue acts, such as ``inform(name = the Wrestlers, price range = cheap, customer rating = low)'') into pictorial representations as used in, e.g. \cite{Williams:2007,black-EtAl:2011:SIGDIAL2011}. We show that pictorial MRs result in better quality NLG data than logic-based textual MRs: utterances elicited by pictorial MRs are judged as significantly more natural, more informative, and better phrased, with a significant increase in average quality ratings (around 0.5 points on a 6-point scale), compared to the logical MRs (see Table~\ref{tab:text-picMR}). Pictorial MRs also result in more spontaneous, natural and varied utterances.
 This is probably due to crowd workers not being primed by lexical tokens. Moreover, as the MR becomes more complex, the benefits of pictorial stimuli increase.

\begin{table}[tp]
\centering
\begin{adjustbox}{max width=0.49\textwidth}
\begin{tabular}{l|l|l}
                & Logic-based MR & Pictorial MR \\
                \hline \hline
informativeness & 4.28**         & \textbf{4.51}**       \\
naturalness     & 4.09**         & \textbf{4.43}**       \\
phrasing        & 4.01**         & \textbf{4.40}**      
\end{tabular}
\end{adjustbox}
\caption{Human evaluation of the data collected with logic-based and pictorial MRs (** denotes $p<$0.01)}
\label{tab:text-picMR}
\end{table}

\section{Conclusion and future work}

Our work addresses two major bottlenecks of current data-driven NLG: Reliable automatic evaluation and efficient high-quality data collection.
While our work shows that we can effectively crowdsource data of sufficient quality to train NLG algorithms -- in particular, using pictorial representations that reduce bias and elicit more syntactically varied and lexically rich data -- our work also clearly demonstrates the need for more advanced evaluation metrics. We see our work as a first step towards reference-less evaluation for NLG by introducing grammar-based metrics.

\section*{Acknowledgements}
\vspace{-0.2cm}
This research received funding from the EPSRC projects  DILiGENt (EP/M005429/1) and  MaDrIgAL (EP/N017536/1). The Titan Xp used for this research was donated by the NVIDIA Corporation.
\bibliography{acl2017}
\bibliographystyle{acl_natbib}

\end{document}